OpenStudy (anonymous):

Les drives 4 miles each way to work at an average of $0.195 per mile. He pays $2.25 a day to park and works an average of 24 days per month. If Les joins a car pool with three other workers, what will his average monthly cost be?

OpenStudy (anonymous):

4 miles each way = 8 miles round trip each day. At .195 per mile (8 x .195) = 1.56 per day (wow...thats cheap gas) He also pays 2.25 per day to park. He works an average of 24 days per month. 24(2.25 + 1.56) = total cost per month.. 54 + 37.44 = $ 91.44 < -- cost per month However, there are 3 other people carpooling with him...which makes a total of 4 people. To figure out what each person pays you need to divide by 4. 91.44/4 = 22.86 (average cost per person) < -- answer Do you have any questions ?

OpenStudy (dwelsher):

22.48------------(1.56*24)(2.25*24)=67.44--------/3= ~22.48